OpenStudy (anonymous):

Could someone please explain this to me in detail? I am struggling with this problem on my homework. K is the midpoint of LM. The coordinates of K are (3, 5) and the coordinates of L are (-5, -6), find the coordinates of M

OpenStudy (jdoe0001):

you recall the mid-point formula, right?

OpenStudy (anonymous):

vaguely.. could you re-explain it to me, please?

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ L&({\color{red}{ -5}}\quad ,&{\color{blue}{ -6}})\quad M&({\color{red}{ M_x}}\quad ,&{\color{blue}{ M_y}}) \end{array}\qquad \\ \quad \\ \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)\implies (3,5) \\ \quad \\ \left(\cfrac{{\color{red}{ M_x}} + {\color{red}{ (-5)}}}{2}\quad ,\quad \cfrac{{\color{blue}{ M_y}} + {\color{blue}{ (-6)}}}{2} \right)\implies (3,5) \begin{cases} \cfrac{{\color{red}{ M_x}} + {\color{red}{ (-5)}}}{2}=3 \\ \quad \\ \cfrac{{\color{blue}{ M_y}} + {\color{blue}{ (-6)}}}{2}=5 \end{cases}\) and solve for the missing coordinate there, in each case, to find the x,y coordinates of point M

OpenStudy (anonymous):

So would it be 11, 16?

OpenStudy (anonymous):

:)

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

thus M ( 11, 16) x y

OpenStudy (anonymous):

Awesome! Could you help me with another, please?

OpenStudy (jdoe0001):

sure... just post anew...thus if I dunno. someone else may know and we can revise each other :)

OpenStudy (anonymous):

OK :)